# Partial domain adaptation enables cross domain cell type annotation between scRNA-seq and snRNA-seq

### v2

@@ 2.5 Shared Source and Target Classifier
The shared source and target classifier operates on the latent representation produced by the encoder to predict the corresponding cell type. It consists of two fully connected layers. The first layer projects the latent features into hidden units and applies a ReLU activation function to enhance non-linear feature interactions. The second layer maps these transformed features to seven output nodes, each representing a distinct cell type. This design allows the classifier to effectively translate the compact latent representation into accurate cell-type predictions while maintaining computational efficiency. The same classifier structure is consistently applied across all experiments without modification.
@@ 2.6 The overview of ScNucAdapt
In conclusion, ScNucAdapt consists of three parts. The framework is shown in Figure 2. The first is the shared encoder for the source and target datasets, aiming to extract representations in the same latent space. The second is the dynamic selection of clusters; in this part, ScNucAdapt is responsible for clustering cell types in target datasets without giving prior knowledge on the number of clusters. Then, we introduce CS Divergence and the rule for merging between the predicted target clusters and the source datasets. ScNucAdapt first trains without merging. After a certain epoch, ScNucAdapt starts calculating CS Divergence and performs merging in terms of the merging rules. Moreover, the total training loss consists of classification loss $L_{cls}$, which is the cross-entropy loss, and the $L_{cs}$ described in section 2.4.
$L=L_{cls}+L_{cs}$ (13)
@@ 2.7 Datasets
Various scRNA-seq and snRNA-seq data are compiled from previous publications. Most datasets are previously preprocessed, the datasets needing preprocessing are preprocessed using scanpy[wolf2018scanpy]. We gathered cells from the bladder, kidney, the mouse cortex, and the frozen and fresh tumor tissues.
For bladder cell types, the dataset is GSE267964[santo2025exploring], which contains two subsets, Immune and Stromal. The datasets are preprocessed in advance and paired.
Moreover, we also collected unpaired scRNA-seq and snRNA-seq of kidney cell types from different publications, GSE140989[menon2020single], which is the scRNA-seq, and GSE121862[lake2019single], which is the snRNA-seq. The cell type labels are gathered from a previous study on annotating cell types in kidneys from scRNA-seq and snRNA-seq using traditional machine learning methods.
@@
For mouse cortical cell types, the datasets are gathered from the GEO database under accession number GSE123454[bakken2018single]. All the cell type labels are collected from previous publications, with each cell annotated.
@@
Table 1: Statistical results of the datasets, including bladder cell types, kidney cell types, frozen and fresh tumors cell types, and mouse cortical cell types
@@ 2.8 Evaluation Metrics
To assess the performance of ScNucAdapt in cell type classification tasks, we evaluated its classification accuracy on datasets with known cell type annotations. The accuracy score quantifies the proportion of correctly predicted cell types among all predictions, providing a straightforward yet informative measure of model performance. Formally, accuracy is defined as shown in Eq 14, where $y_{i}$ denotes the true label of the $i$-th sample, $\hat{y}_{i}$ represents the predicted class label for the same sample, and $n$ is the total number of samples in the dataset. The indicator function $1(y_{i}=\hat{y}_{i})$ returns 1 if the predicted label matches the true label and 0 otherwise.
@@ 2.9 Experimental Setup
The proposed method was evaluated against several existing classifiers that have been widely applied in single-cell transcriptomic analyses. Including SinglecellNet and ScMap. And also a popular domain adaptation method for the cross-batches cell type annotation method in scRNA-seq, ScAdapt[zhou2021scadapt].
In each experiment, we set the learning rate to 0.0001 with the Adam optimizer. We applied early stopping to prevent overfitting. All the experiments are performed on a research server with an NVIDIA GPU GeForce RTX 4090.
@@ 3 Experimental Results
The experimental results shown in Table 3 indicate that ScNucAdapt outperforms existing classifiers, which focus solely on ScRNA-seq datasets. Moreover, ScAdapt. On the immune subset, which is a partial domain adaptation problem where ScRNA-seq is the source data and SnRNA-seq is the target data, ScNucAdapt achieves an accuracy of 90.24, which performed better than ScAdapt’s 90.05, and outperformed SingleCellNet’s 81.02. On the Stromal datasets where ScRNA-seq is the source data and SnRNA-seq is the target data, ScNucAdapt achieves an accuracy of 97.69 under a closed set setting, performing better than ScAdapt’s 96.95, outperforming SingleCellNet’s 91.47. The same holds for the Stromal dataset, where SnRNA-seq serves as the source data and ScRNA-seq as the target data. ScNucAdapt achieves an accuracy score of 90.38, outperforming ScAdapt’s 89.98, SingleCellNet’s 86.61, and ScMap’s 86.97. Interestingly, we found that domain adaptation methods often outperformed scRNA-seq classifiers.
On the unpaired datasets between scRNA-seq and snRNA-seq, which is under a partial setting, ScNucAdapt achieves an accuracy of 87.12, outperforming other methods, including ScAdapt. This might be the cause why ScAdapt tends to negative transfer under partial settings, which is a common situation for traditional domain adaptation methods[cao2018partial]. These results indicate that not only can ScNucAdapt handle distributional differences between scRNA-seq and snRNA-seq, but also under partial settings.
@@
The experimental results presented in Table 4 show that, under a partial adaptation setting for metastatic breast cancer, where ScRNA-seq serves as the source data and SnRNA-seq as the target data, ScAdapt performed slightly better than ScNucAdapt, achieving 94.13, while ScNucAdapt achieved 94.06. Both methods outperform other comparison methods.
Moreover, the cross-domain annotation from scRNA-seq to snRNA-seq on chronic lymphocytic leukemia, ScNucAdapt, achieved an accuracy of 98.39, which performed better than existing scRNA-seq cell type classifiers and ScAdapt. These results demonstrate that ScNucAdapt is effective under partial adaptation settings and can reliably handle cross-domain annotation tasks.
@@ 3.3 ScNucAdapt Enables Cross-Annotation of Mouse Cortical Cell Types Across scRNA-seq and snRNA-seq
The experimental results shown in Table 5 revealed that the cross-annotation from scRNA-seq to snRNA-seq, ScNucAdapt, achieves an accuracy score of 99.78, which performed better than SingleCellNet, ScMap, and scAdapt. On the cross-annotation from snRNA-seq to ScRNA-seq, ScNucAdapt, ScAdapt, and SingleCellNet achieved an accuracy of 100.00, while ScMap achieved 99.13.
@@ 3.5 Sensitive Analysis on C
A total of six cross-domain cell type annotation experiments are included in the analysis. The experimental results shown in Figure 5 that on most occasions, ScNucAdapt is insensitive to the hyperparameters. However, we noticed that there’s a small fluctuation when conducting sensitive analysis on the immune bladder cell type from scRNA-seq to snRNA-seq, but the performance wasn’t significantly degraded.
Overall, the results suggest that ScNucAdapt can automatically adapt to diverse datasets without the need for extensive hyperparameter tuning, thereby improving its applicability in real-world cross-domain annotation between scRNA-seq and snRNA-seq.
@@ 4 Discussion
In this paper, our study fills the research gap on the cross-domain annotation between scRNA-seq and snRNA-seq, and also addresses the distributional and cell composition differences between the two types datasets. The experimental results show the robustness of ScNucAdapt on the cross-domain annotation between scRNA-seq and snRNA-seq, and further ablation experiments have proved the effectiveness of the proposed components in ScNucAdapt. Moreover, insensitive to the hyperparameters that need manual controls on the initial clusters.
@@
First, label noises [chen2025timestamp] that existed in the source datasets could degrade the performance by introducing unreliable supervision signals. These mislabeled samples may hinder domain alignment and reduce the accuracy of downstream annotation.
Second, an exciting direction for future research lies in novel cell type discovery across the two datasets[shi2024scadca]. Current cross-domain annotation frameworks rely heavily on existing cell-type labels and may overlook previously uncharacterized or rare cell populations that are unique to one domain.
Moreover, due to high sparsity and high-dimensional spaces occurring in both scRNA-seq and snRNA-seq, ScNucAdapt tends to overfit, although ScNucAdapt tends to generalize well compared to existing methods on cross-domain annotation between scRNA-seq, there’s still room for improvements on the performance, therefore, future direction could focus on developing algorithms to prevent overfitting and achieve better generalization on target datasets[he2023addressing].
Finally, imbalanced cell type distributions within each domain could hinder generalization. Overrepresented cell types may dominate the training process, causing the model to underperform on rare populations. Future methods could focus on addressing these within-domain imbalances to improve robustness and cross-domain performance [wang2023titok].

### v3

@@ 2.5 Shared Source and Target Classifier
The shared source and target classifier operates on the latent representations produced by the encoder to predict the corresponding cell type. It consists of a single fully connected layer that maps directly to the output nodes, with the number of nodes equal to the number of cell types in the source dataset. This design allows the classifier to effectively translate the compact latent representations into accurate cell-type predictions while maintaining computational efficiency. The same classifier structure is consistently applied across all experiments without modification.
@@ 2.6 The overview of ScNucAdapt
In conclusion, ScNucAdapt consists of three parts. The framework is shown in Figure 3. The first is the shared encoder for the source and target datasets, aiming to extract representations in the same latent space. The second is the dynamic clustering in target representations. In this part, ScNucAdapt is responsible for clustering cell types in target datasets without giving prior knowledge on the number of clusters, and further adjusting through a split and merge framework. Then, we introduce CS Divergence and the rule for merging between the predicted target clusters and the source datasets.
ScNucAdapt employs a two-stage training strategy, and the pseudocode is shown in supplementary S5 Fig. In the first stage, the encoder is trained for T warm-up epochs using only minibatch-based representations learning without clustering, thereby learning meaningful initial feature spaces. In the second stage, each epoch begins by applying GMM clustering and split/merge operation to the full learned representations of the target dataset, followed by source-target matching via argmin of Cauchy-Schwarz divergence between cluster distributions. The encoder is then updated via backpropagation on minibatches using the combined loss shown in Eq. (13) while treating the current cluster assignments as fixed. GMM clustering, split/merge operation, and source-target matching are recomputed every epoch to refine alignments as representations improve. To reduce computational time, we could also recompute every n epochs, but in our experiments, the operation is computed every epoch. Moreover, the total training loss on the selected source and target batch dataset consists of classification loss $L_{cls}$, which is the weighted cross-entropy loss, and the $L_{cs}$ described in section 2.4. $\lambda$ represents the trade-off hyperparameter.
$L=L_{cls}+\lambda*L_{cs}$ (13)
@@ 2.7 Datasets
Various scRNA-seq and snRNA-seq data are compiled from previous publications. Most datasets are preprocessed; datasets that require preprocessing are preprocessed using Scanpy [32]. We gathered cells from the bladder, kidney, the mouse cortex, and the frozen and fresh tumor tissues.
For bladder cell types, the dataset is GSE267964 [25], which contains two subsets, Immune and Stromal. The datasets are preprocessed in advance and paired.
Moreover, we also collected unpaired scRNA-seq and snRNA-seq of kidney cell types from different publications, GSE140989 [18], which is the scRNA-seq, and GSE121862 [10], which is the snRNA-seq. The cell type labels are gathered from a previous study on annotating cell types in kidneys from scRNA-seq and snRNA-seq using traditional machine learning methods.
@@
For mouse cortical cell types, the datasets are gathered from the GEO database under accession number GSE123454 [2]. All the cell type labels are collected from previous publications, with each cell annotated.
@@
Table 1: Statistical results of the datasets, including bladder cell types, kidney cell types, frozen and fresh tumor cell types, and mouse cortical cell types
@@ 2.8 Evaluation Metrics
To assess the performance of ScNucAdapt in cell type classification tasks, we evaluated its classification accuracy on datasets with known cell type annotations. The accuracy score quantifies the proportion of correctly predicted cell types among all predictions, providing a straightforward yet informative measure of model performance. Formally, accuracy is defined as shown in Eq. (14), where $y_{i}$ denotes the true label of the $i$-th sample, $\hat{y}_{i}$ represents the predicted class label for the same sample, and $n$ is the total number of samples in the dataset. The indicator function $1(y_{i}=\hat{y}_{i})$ returns 1 if the predicted label matches the true label and 0 otherwise.
@@
Moreover, we also included the Macro-F1 score to evaluate model performance. Unlike overall accuracy, which can be biased toward majority classes, the Macro-F1 score provides a balanced assessment by treating all classes equally.
First, for each class $c\in\mathcal{C}$ (where $\mathcal{C}$ is the set of all classes), we compute class-specific precision and recall based on the true positives ($TP_{c}$), false positives ($FP_{c}$), and false negatives ($FN_{c}$):
$\text{Precision}_{c}=\frac{TP_{c}}{TP_{c}+FP_{c}}$ (15)
$\text{Recall}_{c}=\frac{TP_{c}}{TP_{c}+FN_{c}}$ (16)
Precision measures the proportion of cells predicted to be class $c$ that are correctly assigned, reflecting the model’s exactness. Recall measures the proportion of true class $c$ cells that were successfully retrieved, reflecting the model’s completeness. The $F1$ score for each class is then defined as the harmonic mean of precision and recall, balancing the trade-off between the two:
$\text{F1}{c}=2\cdot\frac{\text{Precision}{c}\cdot\text{Recall}_{c}}{\text{Precision}_{c}+\text{Recall}_{c}}$ (17)
Finally, the Macro-F1 score is calculated as the arithmetic mean of these per-class F1 scores across all classes:
$\text{Macro-F1}=\frac{1}{|\mathcal{C}|}\sum_{c\in\mathcal{C}}\text{F1}_{c}$ (18)
@@ 2.9 Experimental Setup
The proposed method was evaluated against several existing classifiers that have been widely applied in single-cell transcriptomic analyses. Including SingleCellNet and ScMap. And also a popular domain adaptation method for the cross-batches cell type annotation method in scRNA-seq, ScAdapt [40]. All the comparison methods are tuned to achieve the best performance and are performed on a research server equipped with an NVIDIA GeForce RTX 4090 GPU.
For initial clustering, we applied Gaussian Mixture Models with diagonal covariance matrices set to ’diag’ in scikit-learn, and $C$ as the number of mixture components, which would be further adjusted.
In each experiment, the specific hyperparameters, including the width of the first hiddenlayer, the latent representations width, the early stopping, the trade-off hyperparameter, the learning rate, batchsize, and the epoch before performing initial Gaussian mixture clustering, and the initial cluster number K, are shown in Supplementary S1 Table. Moreover, the optimizer Adam and the hyperparameter $\sigma=5$ are fixed across datasets.
Given a set of counts $\{c_{1},c_{2},\ldots,c_{n}\}$ for $n$ classes, the weight $w_{i}$ for class $i$ is computed as:
$w_{i}=\frac{1}{c_{i}}$ (19)
Then, the weights in cross-entropy loss are calculated as:
$\hat{w}_{i}=\frac{\frac{1}{c_{i}}}{\sum_{j=1}^{n}\frac{1}{c_{j}}}\times n$ (20)
Where $c_{i}$ is the count for class $i$, $n$ is the total number of classes, and $\hat{w}_{i}$ is the final normalized weight for class $i$. Note that we’ve set a slightly higher weight for monocyte and plasma cell in bladder immune scRNA-seq to snRNA-seq. And on CLL, we set a slightly higher weight for the T cell.
The NIW hyperparameters are fixed across experiments and follow the settings from PRAGA and DeepDPM, where we used $\kappa=0.0001$, set $m$ to be the data mean, $\nu$ to be $K+2$, $K$ represents the initial clustering hyperparameter, and $\psi=I\times 0.005$, where $I$ denotes the identity matrix.
@@ 3 Experimental Results
3.1 Simulation Experiments
We first generated simulated datasets using the R package Splatter [37], including both imbalanced and balanced datasets. Each dataset comprises two batches and five cell types. Following this, we performed two target-controlled experiments. In the first, we varied the batch effect strength by adjusting the batch.facScale parameter while keeping the number of cell types in the target dataset fixed at three. In the second experiment, we held the batch.facScale constant at one and varied the number of cell types present in the target data.
Figure 4: Simulation Experiments using Splatter (A) Accuracy on imbalanced and balanced simulated datasets domain shift experiments (B) Macro f1-score on imbalanced and balanced simulated datasets domain shift experiments (C) Accuracy on imbalanced and balanced simulated datasets class variation experiments (D) Macro f1-score on imbalanced and balanced simulated datasets class variation experiments (E) Uncorrected simulated dataset of cell types colored (F) cell type representations before source classes and target clusters merging (G) cell type representations after source classes and target clusters merging (H) Uncorrected simulated dataset of batch colored (I) batch representations before source classes and target clusters merging (J) batch representations after source classes and target clusters merging.
We provide UMAP visualizations shown in Figure 4 to illustrate batch effects and cell type distributions under three conditions: before correction, at the onset of merging, and after complete merging.
The results show that across all conditions and evaluation metrics, ScNucAdapt consistently achieves robust performance. Furthermore, the UMAP visualizations shown in Figure 4 on imbalanced simulated datasets with batch.facScale set to 1, which clearly demonstrates that ScNucAdapt effectively merges the correct cell types while exhibiting minimal negative transfer.
@@ 3.1 Simulation Experiments
The experimental results shown in Table 3 and in Supplementary S2 Table indicate that ScNucAdapt outperforms existing classifiers, which focus solely on scRNA-seq datasets. Moreover, ScAdapt. On the immune subset, which is a partial domain adaptation problem where scRNA-seq is the source data and snRNA-seq is the target data, ScNucAdapt achieves an accuracy of 91.05 and a macro-F1 Score 84.69, which performed better than ScAdapt’s 90.24 (accuracy) and 82.37 (macro-f1), and outperformed SingleCellNet’s 81.02 (accuracy) and 55.51 (macro-f1). On the Stromal datasets where scRNA-seq is the source data, and snRNA-seq is the target data, ScNucAdapt achieves an accuracy of 97.80 and 89.42 of macro-f1 under a closed set setting, performing better than ScAdapt’s 96.95 (accuracy) and 80.00 (macro-f1), outperforming SingleCellNet’s 91.92 (accuracy) and 63.82 (macro-f1). The same holds for the Stromal dataset, where snRNA-seq serves as the source data and scRNA-seq as the target data. ScNucAdapt achieves an accuracy score of 90.38 and 72.42 on macro-f1 score, outperforming ScAdapt’s 89.98 (accuracy) and 71.61 (macro-f1), SingleCellNet’s 86.61 (accuracy) and 66.96 (macro-f1), and ScMap’s 87.02 (accuracy) and 63.72 (macro-f1). Interestingly, we found that domain adaptation methods often outperformed scRNA-seq classifiers. Visualization results using UMAP for the bladder tissue with three subsets are shown in Supplementary S3 Fig. While most scRNA-seq and snRNA-seq populations merged effectively, we observed a limitation in aligning stromal populations, where ScNucAdapt failed to distinguish between vein endothelial cells and general endothelial cells. This is attributable to the extremely limited training set for vein endothelial cells, which contained only four cells. This makes it a challenging scenario for any alignment method. Despite this, ScNucAdapt maintained the best overall performance among all compared methods in terms of both Accuracy and macro F1-score.
On the unpaired datasets between scRNA-seq and snRNA-seq, which is under a partial setting, ScNucAdapt achieves an accuracy of 87.23 and a macro-f1 of 81.5, outperforming other methods, including ScAdapt. The visualization results of UMAP are shown in Figure 5, where most scRNA-seq and snRNA-seq populations are well-merged across batches while maintaining clear separation by cell type. These results indicate that not only can ScNucAdapt handle distributional differences between scRNA-seq and snRNA-seq, but also under partial settings of unpaired scRNA-seq and snRNA-seq.
Figure 5: Visualization result of scRNA-seq and snRNA-seq representations. Using UMAP on kidney tissue (A) visualization result on scRNA-seq to snRNA batch representations (B) visualization result on scRNA-seq to snRNA cell type representations
@@
The experimental results presented in Table 4 and Supplementary S2 Table show that, under a partial adaptation setting for metastatic breast cancer, where scRNA-seq serves as the source data and snRNA-seq as the target data, ScNucAdapt performed better than ScAdapt’s 94.17 (accuracy) and 74.89 (macro-f1), achieving an accuracy of 95.39 and a macro-f1 score of 76.16. Both methods outperform other comparison methods.
Moreover, the cross-domain annotation from scRNA-seq to snRNA-seq on chronic lymphocytic leukemia, ScNucAdapt, achieved an accuracy of 98.39 and a macro-f1 of 94.47, which performed better than existing scRNA-seq cell type classifiers and ScAdapt. These results demonstrate that ScNucAdapt is effective under partial adaptation settings and can reliably handle cross-domain annotation tasks.
Figure 6 presents the UMAP visualization results, illustrating that cells from scRNA-seq and snRNA-seq are well-mixed after integration, yet remain distinctly separated according to their cell type identities.
Figure 6: Visualization result of scRNA-seq and snRNA-seq representations. Using UMAP on frozen and fresh tumor tissue (A) visualization result on scRNA-seq to snRNA batch representations using CLL (B) visualization result on scRNA-seq to snRNA cell type representations using CLL (C) visualization result on scRNA-seq to snRNA batch representations using MBC (D) visualization result on scRNA-seq to snRNA cell type representations using MBC.
@@ 3.4 ScNucAdapt Enables Cross-Annotation of Mouse Cortical Cell Types Across scRNA-seq and snRNA-seq
The experimental results shown in Table 5 and supplementary S3 Table revealed that the cross-annotation from scRNA-seq to snRNA-seq, ScNucAdapt, achieves an accuracy score of 99.78, which performed better than SingleCellNet, ScMap, and scAdapt. On the cross-annotation from snRNA-seq to scRNA-seq, ScNucAdapt, ScAdapt, and SingleCellNet achieved an accuracy of 100.00, while ScMap achieved 99.13 (accuracy) and 97.99 (macro-f1).
As shown in Supplementary S4 Fig, the UMAP embeddings demonstrate successful integration, with cells from both scRNA-seq and snRNA-seq mixing effectively while maintaining clear separation by cell type.
@@ 3.6 Sensitive Analysis
A total of six cross-domain cell type annotation experiments are included in the analysis. The experimental results shown in Figure 8 show that on most occasions, ScNucAdapt is insensitive to the hyperparameters. However, we noticed that there’s a small fluctuation when conducting sensitive analysis on the immune bladder cell type from scRNA-seq to snRNA-seq, but the performance wasn’t significantly degraded. The results suggest that ScNucAdapt can automatically adapt to diverse datasets without the need for extensive hyperparameter tuning, thereby improving its applicability in real-world cross-domain annotation between scRNA-seq and snRNA-seq.
Moreover, we conducted sensitivity analysis on the trade-off hyperparameter $\lambda$ following the six cross-domain cell type annotation experiments. The experimental results shown in Supplementary S1 Fig indicate that ScNucAdapt is not sensitive to the changes of the trade-off hyperparameter across a wide range of values. The model maintained stable performance for $\lambda$, demonstrating that the proposed method achieves consistent domain adaptation without requiring extensive hyperparameter tuning.
3.7 Runtime and Peak GPU Memory Scalability Tests
To assess the runtime and peak GPU memory of ScNucAdapt, we performed scaling experiments on simulated single-cell RNA-seq datasets generated with Splatter. We simulated four datasets with approximately 2,000, 5,000, 10,000, and 20,000 cells (10,000 genes, batch.facScale 0.8), which, after removing two cell types from the target batch, resulted in final sizes of 1,613, 4,013, 7,985, and 16,011 cells, respectively. The results shown in Supplementary S2 Fig show that interestingly, memory consumption scaled linearly from 0.065GB to 0.413GB across these dataset sizes, with peak memory remaining stable during encoder updates due to minibatch-based backpropagation. Runtime per epoch increased from 45.7 seconds to 1,436.4 seconds, with the computational bottleneck being the GMM clustering and split-merge operations performed on the full dataset each epoch. Therefore, we could rerun the clustering and matching every n epochs to reduce runtime.
@@ 4 Discussion
In this paper, our study fills the research gap on the cross-domain annotation between scRNA-seq and snRNA-seq, and also addresses the distributional and cell composition differences between the two types of datasets. The experimental results show the robustness of ScNucAdapt on the cross-domain annotation between scRNA-seq and snRNA-seq, and further ablation experiments have proved the effectiveness of the proposed components in ScNucAdapt. Moreover, insensitive to the hyperparameters that need manual controls on the initial clusters.
@@
First, label noises [3] that existed in the source datasets could degrade the performance by introducing unreliable supervision signals. These mislabeled samples may hinder domain alignment and reduce the accuracy of downstream annotation.
Second, an exciting direction for future research lies in novel cell type discovery across the target dataset [26]. Current cross-domain annotation frameworks rely heavily on existing cell-type labels and may overlook previously uncharacterized or rare cell populations that are domain-specific. Therefore, future work must integrate Open-Set Domain Adaptation [19] or universal domain adaptation frameworks [36], as these are specifically designed to handle both shared and private label sets, allowing models to flag target-domain-specific cells as ”unknown” for further validation.
A more challenging but realistic direction involves scenarios where the gene sets differ substantially between scRNA-seq and snRNA-seq. This moves the problem into the realm of Heterogeneous Domain Adaptation [15], where the feature spaces themselves are mismatched. Future frameworks capable of projecting heterogeneous gene sets into a common latent space would vastly expand the flexibility and applicability of cross-domain annotation.
Moreover, due to high sparsity and high-dimensional spaces occurring in both scRNA-seq and snRNA-seq, ScNucAdapt tends to overfit, although ScNucAdapt tends to generalize well compared to existing methods on cross-domain annotation between scRNA-seq, there’s still room for improvements on the performance. Therefore, future direction could focus on developing algorithms to prevent overfitting and achieve better generalization on target datasets [6].
Finally, imbalanced cell type distributions within each domain could hinder generalization. Overrepresented cell types may dominate the training process, causing the model to underperform on rare populations. Future methods could focus on addressing these within-domain imbalances to improve robustness and cross-domain performance [31].